# Doomed Worlds I: No new evidence for orbital decay in a long-term survey of 43 ultra-hot Jupiters

### v1

@@ Abstract
Ultra-hot Jupiters are likely doomed by tidal forces to undergo orbital decay and eventual disruption by their stars, but the timescale over which this process unfolds is unknown. We present results from a long-term project to monitor ultra-hot Jupiters transits. We recovered WASP-12 b’s orbital decay rate of ms yr, in agreement with prior work. Five other systems initially had promising non-linear transit ephemerides. However, a closer examination of two – WASP-19 b and CoRoT-2 b, both with prior tentative detections – revealed several independent errors with the literature timing data; after correction neither planet shows signs of orbital decay. Meanwhile, a potential decreasing period for TrES-1 b, ms yr, corresponds to a tidal quality factor and likely does not result from orbital decay, if driven by dissipation within the host star. Nominal period increases in two systems, WASP-121 b and WASP-46 b, rest on a small handful of points. Only 1/43 planets (WASP-12 b) in our sample is experiencing detectable orbital decay. For nearly half (20/42) we can rule out as high as observed for WASP-12 b. Thus while many ultra-hot Jupiters could still be experiencing rapid decay that we cannot yet detect, a sizeable sub-population of UHJs are decaying at least an order of magnitude more slowly than WASP-12 b. Our reanalysis of Kepler-1658 b with no new data finds that it remains a promising orbital decay candidate. Finally, we recommend that the scientific community take steps to avoid spurious detections through better management of the multi-decade-spanning datasets needed to search for and study planetary orbital decay.
@@ 1 Introduction
Ultra-hot Jupiters (UHJs), or giant planets with orbital periods less than about 3 days, experience significant tidal effects that play a critical role in their long-term dynamical stability. Recent work has shown that the population of all hot Jupiter ( d) host stars is younger than the general population of either field stars or planet-hosting stars (Hamer & Schlaufman, 2019), and two separate analyses have found lower UHJ occurrence rates around older host stars (Miyazaki & Masuda, 2023; Chen et al., 2023). All of this supports the idea that many giant, close-in planets quickly inspiral or are otherwise destroyed while their stars are still on the main sequence. Meanwhile, an estimated half of all stars may have ingested a former UHJ in the first Gyr of the star’s lifetime (Matsakos & Königl, 2015). Recent direct detection of a planetary engulfment by De et al. (2023) for an inferred planet of 0.1-10 came in the form of a low-luminosity optical transient lasting several days, followed by infrared brightening for several months. Such planet engulfment events are thought to occur somewhere in the galaxy about once every 1-10 years (Metzger et al., 2012; De et al., 2023). Estimates for the lifetimes of UHJs as a population have been limited, however, by theoretical uncertainties in the value of the stellar tidal dissipation factor, , where estimates range from (Jackson et al., 2008; Husnoo et al., 2012) to (Penev et al., 2012; Collier Cameron & Jardine, 2018). More recent theoretical work that incorporates nonlinear dissipation has found important variations in with stellar mass and age and that planets around stars with only experience significant tidal decay when the star is on the subgiant branch (Weinberg et al., 2024). Penev et al. (2018) found a steep dependence on forcing frequency for , with for orbital periods and increasing rapidly to for . Consequently, tidal decay might be expected to slow as the planet nears its star. It is important to note, however, that for planets whose rotation states are not tidally locked or whose orbits are eccentric, tidal decay can be driven by dissipation within the planet. However, tidal locking (Guillot & Showman, 2002) and orbital circularization (Dawson & Johnson, 2018) timescales for hot Jupiters are thought to be less than millions of years. Thus, in the absence of an exotic rotation state (Millholland & Laughlin, 2018; Efroimsky & Makarov, 2022) or dynamical excitation of eccentricity (Pu & Lai, 2019), the contribution to tidal decay from dissipation within the planets is likely to be short-lived.
@@
With a known population of around 100 UHJs, it is now possible to empirically address the open questions of how many UHJs are in decaying orbits and at what rates they are decaying. Strong evidence for decay has been presented for just two UHJs, WASP-12 b (Patra et al., 2017; Baluev et al., 2019; Yee et al., 2020) and Kepler-1658 b (Vissapragada et al., 2022), with orbital periods that are apparently decreasing by ms yr and ms yr, respectively, corresponding to remaining lifetimes of 3 Myr and 2.5 Myr. These decay rates also correspond to stellar tidal dissipation parameters of and , respectively. Kepler-1658 b orbits an evolved subgiant (, ), for which that rapid rate of decay matches the theoretical predictions of Vissapragada et al. (2022), though perhaps not those of Barker et al. (2024) (see discussion in Section 6.7). Few planets are known around recently evolved stars due to challenges with their detection (see e.g. Johnson et al., 2007). In fact, one explanation for the evolution of WASP-12 b may be that its host star is also a subgiant (Weinberg et al., 2017), though modeling and observational uncertainties leave its status unclear. This raises an open question: are there any ultra-hot Jupiters around main sequence stars with decaying orbits?
@@
In the past few years, as observational baselines have passed the decade mark for many systems, a growing number of UHJs around main-sequence stars have been presented with suggestions that they may have decreasing orbital periods. These include: HAT-P-19 b (Hagey et al., 2022); HAT-P-32 b (Hagey et al., 2022); HAT-P-51b (Yeh et al., 2023); HAT-P-53b (Yeh et al., 2023); KELT-9 b (Harre et al., 2023); TrES-1 b (Hagey et al., 2022; Ivshina & Winn, 2022); TrES-2 b (Hagey et al., 2022); TrES-3 b (but noted to be marginal, Mannaday et al., 2022; Hagey et al., 2022); TrES-5 b (Maciejewski et al., 2021; Ivshina & Winn, 2022; Yeh et al., 2023); WASP-4b (Bouma et al., 2020; Harre et al., 2023); WASP-19 b (Patra et al., 2020; Ivshina & Winn, 2022); WASP-32 b (but not significant, Sun et al., 2023); WASP-43 b as reported by Sun et al. (2018) though not by Hagey et al. (2022); and XO-3 b (Yang & Wei, 2022; Ivshina & Winn, 2022). Many of these claimed detections are acknowledged in the original publications to be of marginal significance, however. It is also common for an apparent non-linear ephemeris based on a small numbers of transits and/or a short baseline to disappear with additional observation (e.g., as for OGLE-TR-113b, Adams et al., 2010; Hoyer et al., 2016). Even after an ostensible period decay is detected, it is necessary to rule out other physical mechanisms, such as apsidal precession (the current best explanation for KELT-9b, Harre et al., 2023), perturbations caused by companion planets or stars, or even the acceleration of the host star towards the Earth (proposed as the explanation for WASP-4b, Bouma et al., 2020).
@@
Note that while the parameter used in the fit is typically (which has units of days per epoch), the value most frequently quoted is the rate of change in ms per year, or
where P is the orbital period in days, is in days per epoch, and is the rate of change of P in ms yr.
@@
In this work, we present the first results of a new, long-term program to take regular ground-based observations of UHJs, and combine these data with times from the literature, which for some planets stretch back over two decades. We have observed over a hundred new light curves of 43 UHJs for this project since 2022, and we are also publishing a cache of older light curves from 2008-2009 that had previously eluded publication. For this first paper, we summarize the results for all 43 systems but leave an in-depth analyses of most systems for future work. We focus here on six planets – CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b – all of which initially had values for from calculations using our new midtimes and the midtimes recorded in the literature.
It is important to state from the outset that for two of the six planets we chose to focus on in this work – CoRoT-2 b and WASP-19 b, which initially had the largest values after WASP-12 b – we needed to correct numerous errors and inconsistencies in the literature. After doing so, our final values for both planets dropped so significantly that neither one of them now shows any signs of non-linear ephemerides. WASP-12 b remains the only candidate in our sample with evidence for orbital decay. (Notably, our sample does not include Kepler-1658 b, for which we have no new observations: its depth of 0.1% makes it difficult to observe except from space or from the largest ground-based observatories. We do however apply the same analysis to the existing data for Kepler-1658 b and find it to be a very promising candidate, as discussed in Section 6.7.)
@@
| Planet | Date | Instr. | Time | Filter | Exp. | Aper. | Scatter | |||
@@
| WASP-12b | 2022-10-28 | IoIO | 4.3 | 179 | 3 | 8 | R | 42.71 | NA | 0.32 |
| WASP-12b | 2022-11-09 | IoIO | 4.4 | 276 | 5 | 1 | R | 42.71 | NA | 0.55 |
| WASP-12b | 2023-01-30 | IoIO | 4.9 | 329 | 0 | 12 | R | 42.71 | NA | 0.4 |
| WASP-12b | 2023-02-11 | IoIO | 4.1 | 221 | 28 | 1 | R | 42.71 | NA | 1.11 |
| WASP-12b | 2023-03-05 | IoIO | 3.6 | 237 | 2 | 8 | R | 42.71 | NA | 0.41 |
| WASP-12b | 2023-04-22 | IoIO | 3.6 | 202 | 14 | 16 | R | 42.71 | NA | 0.82 |
@@ 2 New Ground-based Observations and Photometry
In this paper, we are publishing the photometry for twenty-three new transits of six planets: CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b, shown in Figure 1. (To allow for more in-depth analyses of individual systems, transit light curve photometry and full timing analyses for the other 37 planets in our sample will be published in future works. The results presented in this work for those 37 systems are based on at least one new light curve along with the literature transit midtimes compiled as in Section 5, although no attempt has been to correct literature times except for removing obvious duplicates.) Five ground-based observatories produced these observations in 2008-2009 and 2022-2023, with observational details shown in Table 1. All ground-based, TESS, and CoRoT light curves for the six planets featured in this work are available as supplementary data, with a stub table for format in Table 2.
| Transit | Planet | Instrument | Time (BJD/TDB) | Flux | Model |
@@ 2.1 IoIO observations
The Planetary Science Institute’s Io Input/Output Observatory (IoIO) is a small-aperture (35-cm) robotic telescope located at the San Pedro Valley Observatory, a hosting site situated in a dark location 100 km east of Tucson, Arizona. IoIO was purpose-built to record observations of faint gases around Jupiter that trace their origin from Jupiter’s moon Io, and as such has a built-in coronagraph (for more details about the observatory, see e.g. Morgenthaler et al., 2019), but it is also a fully functional general-purpose telescope. Since 2022, much of the time when Jupiter is not available has been dedicated to exoplanet transits, and we have successfully observed over 100 full or partial exoplanet transits, six of whose light curves are published in this work. Observations in 2022-2023 were scripted to be taken in the R filter at an exposure time chosen for the magnitude of the host star. For two transits of WASP-12 b, bad weather or scheduling issues resulted in only partial observations over egress.
@@ 2.2 SAAO 40-inch observations
Observations were taken on the 40-inch telescope operated by the South African Astronomical Observatory (SAAO) in Sutherland, South Africa, using the Sutherland High-speed Optical Camera (SHOC) facility instruments (Coppejans et al., 2013). These instruments are ideal for exoplanet observations, as they were built to be optimized for high-speed, accurately-timed imaging. They employ Andor iXon 888 frame-transfer EMCCD cameras which are triggered by GPS, mounted with a suite of Johnson-Cousins and Sloan filters, as well as an open-filter setting. Observations were mostly taken in , with other details in Table 1. All data were taken with 2x2 binning (0.334 arcsec per superpixel), the 3 MHz conventional mode, and 5.2x gain setting (8.6 electrons/ADU). Biases and twilight flats were taken each night of the observations, and all data were bias-subtracted and flatfielded. For two transits of CoRoT-2 b and one transit of WASP-12 b, bad weather and/or scheduling issues resulted in only partial observations over ingress and/or egress.
@@ 2.3 Differential photometry of IoIO and SAAO 40-inch data
For each target star, the pipeline specifies a default range of star brightnesses relative to the target for acceptable comparison stars. It then calculates the ratio of the target star to each potential companion star, and rejects those with scatter above a particular threshold (adjustable for each target star). Comparison stars that were not present on every field (typically due to tracking motion or to variable cloud conditions causing stars to either saturate or become too faint) are automatically removed. Other problematic stars, e.g. strongly variable stars, can be excluded for individual transits as needed. There is also a mechanism to identify bad frames in cases where all comparison stars are lost; after omitting those frames, comparison star selection proceeds as above using the remaining good frames. Table 1 shows the number of frames and comparison stars used for each light curve. Once the best set of comparison stars has been identified, a normalized transit light curve is produced for each aperture: the counts from the target star are divided by the sum of the counts of the companion stars, then normalized by dividing by the median flux outside of the predicted transit window. The best aperture is the one that minimizes the scatter on the out-of-transit baseline (see Table 1). Sometimes the resulting transit has a slope in the overall flux, often correlated with the airmass of the target star. We leave detrending any such slopes as a free parameter during fitting (Section 4).

### v2

@@ Abstract
Ultrahot Jupiters (UHJs) are likely doomed by tidal forces to undergo orbital decay and eventual disruption by their stars, but the timescale over which this process unfolds is unknown. We present results from a long-term project to monitor UHJ transits. We recovered WASP-12 b’s orbital decay rate of ms yr-1, in agreement with prior work. Five other systems initially had promising nonlinear transit ephemerides. However, a closer examination of two – WASP-19 b and CoRoT-2 b, both with prior tentative detections – revealed several independent errors with the literature timing data; after correction neither planet shows signs of orbital decay. Meanwhile, a potential decreasing period for TrES-1 b, ms yr-1, corresponds to a tidal quality factor and likely does not result from orbital decay, if driven by dissipation within the host star. Nominal period increases in two systems, WASP-121 b and WASP-46 b, rest on a small handful of points. Only 1/43 planets (WASP-12 b) in our sample is experiencing detectable orbital decay. For nearly half (20/42) we can rule out as high as observed for WASP-12 b. Thus while many ultra-hot Jupiters could still be experiencing rapid decay that we cannot yet detect, a sizable sub-population of UHJs are decaying at least an order of magnitude more slowly than WASP-12 b. Our reanalysis of Kepler-1658 b with no new data finds that it remains a promising orbital decay candidate. Finally, we recommend that the scientific community take steps to avoid spurious detections through better management of the multi-decade-spanning datasets needed to search for and study planetary orbital decay.
@@ 1 Introduction
Ultrahot Jupiters (UHJs), or giant planets with orbital periods less than about 3 days, experience significant tidal effects that play a critical role in their long-term dynamical stability. Recent work has shown that the population of all hot Jupiter ( d) host stars is younger than the general population of either field stars or planet-hosting stars (Hamer & Schlaufman, 2019), and two separate analyses have found lower UHJ occurrence rates around older host stars (Chen et al., 2023; Miyazaki & Masuda, 2023). All of this supports the idea that many giant, close-in planets quickly inspiral or are otherwise destroyed while their stars are still on the main sequence. Meanwhile, an estimated half of all stars may have ingested a former UHJ in the first Gyr of the star’s lifetime (Matsakos & Königl, 2015). Recent direct detection of a planetary engulfment by De et al. (2023) for an inferred planet of 0.1-10 came in the form of a low-luminosity optical transient lasting several days, followed by infrared brightening for several months. Such planet engulfment events are thought to occur somewhere in the galaxy about once every 1-10 years (Metzger et al., 2012; De et al., 2023). Estimates for the lifetimes of UHJs as a population have been limited, however, by theoretical uncertainties in the value of the stellar tidal dissipation factor, , where estimates range from (Jackson et al., 2008; Husnoo et al., 2012) to (Penev et al., 2012; Collier Cameron & Jardine, 2018). More recent theoretical work that incorporates nonlinear dissipation has found important variations in with stellar mass and age and that planets around stars with only experience significant tidal decay when the star is on the subgiant branch (Weinberg et al., 2024). Penev et al. (2018) found a steep dependence on forcing frequency for , with for orbital periods and increasing rapidly to for . Consequently, tidal decay might be expected to slow as the planet nears its star. It is important to note, however, that for planets whose rotation states are not tidally locked or whose orbits are eccentric, tidal decay can be driven by dissipation within the planet. However, tidal locking (Guillot & Showman, 2002) and orbital circularization (Dawson & Johnson, 2018) timescales for hot Jupiters are thought to be less than millions of years. Thus, in the absence of an exotic rotation state (Millholland & Laughlin, 2018; Efroimsky & Makarov, 2022) or dynamical excitation of eccentricity (Pu & Lai, 2019), the contribution to tidal decay from dissipation within the planets is likely to be short-lived.
@@
With a known population of around 100 UHJs, it is now possible to empirically address the open questions of how many UHJs are in decaying orbits and at what rates they are decaying. Strong evidence for decay has been presented for just two UHJs, WASP-12 b (Patra et al., 2017; Baluev et al., 2019; Yee et al., 2020) and Kepler-1658 b (Vissapragada et al., 2022), with orbital periods that are apparently decreasing by ms yr-1 and ms yr-1, respectively, corresponding to remaining lifetimes of 3 Myr and 2.5 Myr. These decay rates also correspond to stellar tidal dissipation parameters of and , respectively. Kepler-1658 b orbits an evolved subgiant (, ), for which that rapid rate of decay matches the theoretical predictions of Vissapragada et al. (2022), though perhaps not those of Barker et al. (2024) (see discussion in Section 6.7). Few planets are known around recently evolved stars due to challenges with their detection (see e.g. Johnson et al., 2007). In fact, one explanation for the evolution of WASP-12 b may be that its host star is also a subgiant (Weinberg et al., 2017), though modeling and observational uncertainties leave its status ambiguous (Leonardi et al., 2024). This raises an open question: are there any ultra-hot Jupiters around main-sequence stars with decaying orbits?
@@
In the past few years, as observational baselines have passed the decade mark for many systems, a growing number of UHJs around main-sequence stars have been presented with suggestions that they may have decreasing orbital periods. These include: HAT-P-19 b (Hagey et al., 2022); HAT-P-32 b (Hagey et al., 2022); HAT-P-51 b (Yeh et al., 2024); HAT-P-53 b (Yeh et al., 2024); KELT-9 b (Harre et al., 2023); TrES-1 b (Hagey et al., 2022; Ivshina & Winn, 2022); TrES-2 b (Hagey et al., 2022); TrES-3 b (but noted to be marginal, Mannaday et al., 2022; Hagey et al., 2022); TrES-5 b (nonlinearity though not decay noted by Maciejewski et al., 2021; Hagey et al., 2022; Ivshina & Winn, 2022; Yeh et al., 2024); WASP-4 b (Bouma et al., 2020; Hagey et al., 2022; Harre et al., 2023); WASP-19 b (Patra et al., 2020; Ivshina & Winn, 2022); WASP-32 b (but not significant, Sun et al., 2023); WASP-43 b as reported by Sun et al. (2018) though not by Hagey et al. (2022); and XO-3 b (Yang & Wei, 2022; Ivshina & Winn, 2022). Many of these claimed detections are acknowledged in the original publications to be of marginal significance, however. It is also common for an apparent non-linear ephemeris based on a small numbers of transits and/or a short baseline to disappear with additional observation (e.g., as for OGLE-TR-113b, Adams et al., 2010; Hoyer et al., 2016). Even after an ostensible period decay is detected, it is necessary to rule out other physical mechanisms, such as apsidal precession (the current best explanation for KELT-9 b, Harre et al., 2023), perturbations caused by companion planets or stars, or even the acceleration of the host star towards the Earth (proposed as the explanation for WASP-4 b, Bouma et al., 2020).
@@
Note that while the parameter used in the fit is typically (which has units of days per epoch), the value most frequently quoted in the literature is the rate of change in ms yr-1, or
where is the orbital period in days, is in days per epoch, and is the rate of change of in ms yr-1.
@@
In this work, we present the first results of a new, long-term program to take regular ground-based observations of UHJs and combine these data with times from the literature, which for some planets stretch back over two decades. We have observed more thana hundred new light curves of 43 UHJs for this project since 2022, and we are also publishing a cache of older light curves from 2008-2009 that had previously eluded publication. For this first paper, we summarize the results for all 43 systems but leave an in-depth analysis of most systems for future work. We focus here on six planets – CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b – all of which initially had values for from calculations using our new midtimes and the midtimes recorded in the literature.
It is important to state from the outset that for two of the six planets we chose to focus on in this work – CoRoT-2 b and WASP-19 b, which initially had the largest values after WASP-12 b – we needed to correct numerous errors and inconsistencies in the literature. After doing so, our final values for both planets dropped so significantly that neither one of them now shows any signs of nonlinear ephemerides. WASP-12 b remains the only candidate in our sample with evidence for orbital decay. (Notably, our sample does not include Kepler-1658 b, for which we have no new observations: its depth of 0.1% makes it difficult to observe except from space or from the largest ground-based observatories. We do however apply the same analysis to the existing data for Kepler-1658 b and find it to be a very promising candidate, as discussed in Section 6.7.)
@@
| Planet | Datea | Instr. | Time | b | Filter | Exp. | Aper. | Scatterc | ||
@@
| WASP-12b | 2022-10-28 | IoIO | 4.3 | 179 | 3 | 8 | R | 42.71 | NAd | 0.32 |
| WASP-12b | 2022-11-09 | IoIO | 4.4 | 276 | 5 | 1 | R | 42.71 | NAd | 0.55 |
| WASP-12b | 2023-01-30 | IoIO | 4.9 | 329 | 0 | 12 | R | 42.71 | NAd | 0.4 |
| WASP-12b | 2023-02-11 | IoIO | 4.1 | 221 | 28 | 1 | R | 42.71 | NAd | 1.11 |
| WASP-12b | 2023-03-05 | IoIO | 3.6 | 237 | 2 | 8 | R | 42.71 | NAd | 0.41 |
| WASP-12b | 2023-04-22 | IoIO | 3.6 | 202 | 14 | 16 | R | 42.71 | NAd | 0.82 |
@@ 2 New Ground-based Observations and Photometry
In this paper, we are publishing the photometry for twenty-three new transits of six planets: CoRoT-2 b, TrES-1 b, WASP-12 b, WASP-19 b, WASP-46 b, and WASP-121 b, shown in Figure 1. (To allow for more in-depth analyses of individual systems, transit light curve photometry and full timing analyses for the other 37 planets in our sample will be published in future works. The results presented in this work for those 37 systems are based on at least one new light curve along with the literature transit midtimes compiled as in Section 5, although no attempt has been made to correct literature times except for removing obvious duplicates.) Five ground-based observatories produced these observations in 2008-2009 and 2022-2023, with observational details shown in Table 1. All ground-based, TESS, and CoRoT light curves for the six planets featured in this work are available as supplementary data, with a stub table for format in Table 2.
| Transit | Planet | Instrumentc | Time (BJD/TDB)d | Fluxe | Modelf |
@@ 2.1 IoIO observations
The Planetary Science Institute’s Io Input/Output Observatory (IoIO) is a small-aperture (35 cm) robotic telescope located at the San Pedro Valley Observatory, a hosting site situated in a dark location 100 km east of Tucson, Arizona. IoIO was purpose-built to record observations of faint gases around Jupiter that trace their origin from Jupiter’s moon Io and as such has a built-in coronagraph (for more details about the observatory, see e.g. Morgenthaler et al., 2019), but it is also a fully functional general-purpose telescope. Since 2022, much of the time when Jupiter is not available has been dedicated to exoplanet transits, and we have successfully observed over 100 full or partial exoplanet transits, six of whose light curves are published in this work. Observations in 2022-2023 were scripted to be taken in the R filter at an exposure time chosen for the magnitude of the host star. For two transits of WASP-12 b, bad weather or scheduling issues resulted in only partial observations over egress.
@@ 2.2 SAAO 40-inch observations
Observations were taken on the 40-inch telescope operated by the South African Astronomical Observatory (SAAO) in Sutherland, South Africa, using the Sutherland High-speed Optical Camera (SHOC) facility instruments (Coppejans et al., 2013). These instruments are ideal for exoplanet observations, as they were built to be optimized for high-speed, accurately timed imaging. They employ Andor iXon 888 frame-transfer EMCCD cameras that are triggered by GPS, mounted with a suite of Johnson-Cousins and Sloan filters, as well as an open-filter setting. Observations were mostly taken in , with other details in Table 1. All data were taken with 2x2 binning (0.334 arcsec per superpixel), the 3 MHz conventional mode, and 5.2x gain setting (8.6 electrons/ADU). Biases and twilight flats were taken each night of the observations, and all data were bias-subtracted and flatfielded. For two transits of CoRoT-2 b and one transit of WASP-12 b, bad weather and/or scheduling issues resulted in only partial observations over ingress and/or egress.
@@ 2.3 Differential photometry of IoIO and SAAO 40-inch data
For each target star, the pipeline specifies a default range of star brightnesses relative to the target for acceptable comparison stars. It then calculates the ratio of the target star to each potential companion star and rejects those with scatter above a particular threshold (adjustable for each target star). Comparison stars that were not present on every field (typically due to tracking motion or to variable cloud conditions causing stars to either saturate or become too faint) are automatically removed. Other problematic stars, e.g. strongly variable stars, can be excluded for individual transits as needed. There is also a mechanism to identify bad frames in cases where all comparison stars are lost; after omitting those frames, comparison star selection proceeds as above using the remaining good frames. Table 1 shows the number of frames and comparison stars used for each light curve. Once the best set of comparison stars has been identified, a normalized transit light curve is produced for each aperture: the counts from the target star are divided by the sum of the counts of the companion stars, then normalized by dividing by the median flux outside of the predicted transit window. The best aperture is the one that minimizes the scatter on the out-of-transit baseline (see Table 1). Sometimes the resulting transit has a slope in the overall flux, often correlated with the airmass of the target star. We leave detrending any such slopes as a free parameter during fitting (Section 4).